OpenStudy (anonymous):

implify the following expression: (x + 6y) - (3x − 10y). If the final answer is written in the form Ax + By, what is the value of A? Numerical Answers Expected

OpenStudy (kinggeorge):

\[(x+6y)-(3x-10y)=x+6y-3x+10=-2x+16y\]From here, it is easy to see that \(A=-2\)

OpenStudy (anonymous):

Thank you.

OpenStudy (kinggeorge):

You're welcome.